What is the pH of an aqueous solution made by combining 43.99 mL of a 0.4070 M sodium formate with 44.43 mL of a 0.3896 M solution of formic acid?
Answer #1

Total volume = 43.99ml + 44.43ml = 88.42ml

Dilution of Formic acid = 88.42ml /44.43ml = 1.99 time

Final [ HCOOH ] = 0.3896M/1.99 =0.1958M

Dilution of Formate = 88.42ml/43.99ml = 2.01time

Final [ HCOO-] = 0.4070M/2.01= 0.2025M

Henderson - Hasselbalch equation is

pH = pKa + log([A-]/[HA])

pKa of Formic acid = 3.75

pH = 3.75 + log (0.2025M/0.1958M)

= 3.75 + 0.01

= 3.76